OpenStudy (ihelpyoulearn):

Simplify (2sqrt5+3sqrt7)^2. I got 83+12sqrt35. Please show how you solved this and if I'm correct. Thanks everyone! Will medal :)

OpenStudy (johnweldon1993):

\[\large (2\sqrt{5} + 3\sqrt{7})^2\] \[\large (2\sqrt{5} + 3\sqrt{7})(2\sqrt{5} + 3\sqrt{7})\] \[\large (4\times 5) + (6\sqrt{35}) + (6\sqrt{35}) + (9\times 7)\] \[\large 83 + 12\sqrt{35}\] \(\large \color \red{\checkmark}\) Good job! :)

OpenStudy (ihelpyoulearn):

Yay! Thanks! @johnweldon1993

OpenStudy (johnweldon1993):

No problem!